OpenStudy (anonymous):

How many grams of chlorine gas is required to completely react with 0.400g of hydrogen gas to yield hydrogen chloride gas

OpenStudy (abmon98):

H2+Cl2-->2HCL Number of Moles=Mass(g)/Molar Mass(g/mol) 0.400/1=0.400 moles Mole Ratio of H2 and Cl2 is 1:1 so they share the same number of moles Arrange the equation: Mass(g)=Molar Mass(g/mol)*Number of Moles 0.400*35.5=14.2

OpenStudy (anonymous):

Also calculate the amount of HCl formed?

OpenStudy (abmon98):

Why dont you try solving it

OpenStudy (anonymous):

cause I don't know..pls

OpenStudy (abmon98):

Look at the mole ratio from the balanced chemical equation of H2 Cl2 and HCL 1:1:2 0.4:0.4:0.8 Mass(g)=Molar Mass(g/mol)*Number of moles 0.8*(1+35.5)=Mass

OpenStudy (anonymous):

thanks :)

OpenStudy (abmon98):

Your welcome :)